OpenStudy (anonymous):

Rectangle EFGH has the following side measures. EF= 54in FG= 891in GH= (1/2x - 3y) in HE= (x + 54y) in What are the values of x and y ?

OpenStudy (anonymous):

x+54y=891 (1/2x-3y)=54 solve for x and y